# Exploring the topological sector optimization on quantum computers

### v2

@@ VI Discussions and Conclusions
On the contrary, QITE algorithm and its NISQ approximate implementation can both successfully address TSO problems. Not only QITE has quadratic advantage on time complexity over QA, but its mechanism which utilizes the property of quantum superposition prevents it from being trapped in some local minima. Topological obstructions have no effects on QITE and thus QITE is a better algorithm for tackling TSO problem than QA. We have to emphasize that the two TSO problems we consider in this work do not even introduce external magnetic fields and it is yet difficult for QA and SQA. If some small random external fields are added, the energy landscape in each sector will be much more glassy and the problems will be more difficult. Unquestionably, QITE can still find the solution as we discussed. The fate of VQITE and VQE in this case is an interesting problem to explore in the future.
@@ Acknowledgements.
We wish to thank Xiaopeng Li and Shuai Yin for helpful discussions. Y.M.D. and Z.Y. thank the start-up fund of Westlake University. The authors acknowledge Beijng PARATERA Tech Co.,Ltd. (https://www.paratera.com/) for providing HPC resources that have contributed to the research results reported within this paper. Y.C.W. acknowledges support from Zhejiang Provincial Natural Science Foundation of China (Grant Nos. LZ23A040003).
Appendix A Simulation of QA
For the QA Hamiltonian $\hat{H}(s)$, the immediate quantum state $|\Psi(t)\rangle$ and the state evolved after a small time interval $\Delta t$ are given by
$|\Psi(t+\Delta t)\rangle=e^{-i\hat{H}(s)\Delta t}|\Psi(t)\rangle,$ (17)
where $s$ is changed by $\Delta s$ every $\Delta t$, thus $\Delta s$ should be small enough compared with $\Delta t$ for the purpose of adiabatic approximation. The total number of evolution steps is therefore $N=1/\Delta t$ and the total evolution time is $T=N\Delta t$. The ground state of $\hat{H}_{1}$ is in (3) is chosen to be $|\Psi(0)\rangle=\otimes_{j}|-\rangle_{j}$, where $|-\rangle=(|0\rangle-|1\rangle)/\sqrt{2}$. We fix $\Delta t=0.1$ and modify $\Delta s$ to get different $T$.
As an example, we test a 16-qubit AFM Ising chain (PBC), whose Hamiltonian is
It turns out that $\Delta s=10^{-4}$, i.e. $T=10^{3}$, is sufficient for this model to reach its ground state $E_{0}$ via QA (See Fig. 10).
Figure 10: Real-time simulation of QA for a 16-qubit AFM Ising chain under periodic boundary condition.
Appendix B Simulation of SQA
@@
For $\hat{H}_{1}$ in the QA Hamiltonian (1), QA always applys uniform magnetic fields on all sites. Different from QA, the SQA method introduces virtual edges to try overcoming the topological obstructions, where an edge along a specified direction will be opened via increasing the intensity of the magnetic fields that acts on it and then will be glued by decreasing the intensity [11].
@@
Here we take the $\hat{H}_{\text{sq}}$ on a $4\times 4$ lattice as an example. There are four edges along the horizontal axis (See Fig. 11). For convenience, we denote $s=t/T$. We use $\mathfrak{h}_{j}(s)$ to represent the intensities of magnetic fields on the $j$th edge and $\overline{\mathfrak{h}}_{j}(s)$ to represent the intensities on those sites that are not included in the $j$th edge at time $s$ or $t=sT$. Furthermore, $s_{\text{max}}<1$ is the time that we regard all the edges have been glued well and $h_{\text{max}}>1$ is the maximum value of $\mathfrak{h}_{j}(s)$.
@@
As it shown in Fig. 11, $\mathfrak{h}_{1}(0)=h_{\text{max}}$ means the first edge has been opened at the beginning. The gluing process of the first edge corresponds to $\mathfrak{h}_{1}(s)=-4h_{\text{max}}s/s_{\text{max}}+h_{\text{max}}$ when $\overline{\mathfrak{h}}_{1}(s)=1-s$, where $s\in[0,(1-h_{\text{max}})/(1-4h_{\text{max}}/s_{\text{max}})]$. Then, a standard QA will be appiled to all sites until $s=s_{\text{max}}/4$. After that, we open the second edge by increasing $\mathfrak{h}_{2}(s_{\text{max}}/4)$ from $41-s_{\text{max}}/$ 4 to $h_{\text{max}}$. Attention that the standard QA is halt for the moment until $\mathfrak{h}_{2}(s_{\text{max}})=h_{\text{max}}$. Next, we similarly glue the second edge as what we do to the first edge. When $s=-4h_{\text{max}}s/s_{\text{max}}+4h_{\text{max}}$, a standard QA will be performed until $s=1$.
@@
Considering the time $T_{\text{open}}$ for opening the edges, the total evolution time should be $T=T_{\text{evo}}+T_{\text{open}}$, where $T_{\text{evo}}$ stands for the time when $s$ is changed. In our simulations for both $\hat{H}_{\text{tri}}$ and $\hat{H}_{\text{sq}}$, we take $h_{\text{max}}=2$ and $s_{\text{max}}=0.8$, which are same as that in Ref. [11]. Besides, the choice of $\Delta t$ and $\Delta s$ follows Appx. A.
@@
Appendix C Details of VQITE, Diag-VQITE and VQE
@@
where $H$ denotes Hadamard gate. The variational parameters of the PQC are initially set to follow Gaussian distribution ($\mu=0,\sigma=0.05$). For updating the parameters, we add a small quantity $\epsilon=10^{-4}$ to all the digonal elements $A_{jj}$ in Eq. (11) for the purpose of regularization when solving the linear equations in VQITE and Diag-VQITE. In VQE, we use the stochastic gradient descent to minimize $\langle\phi(\vec{\theta})|\hat{H}|\phi(\vec{\theta})\rangle$.

### v3

@@ VI Discussions and Conclusions
It is worth of emphasizing that for some TSO models, it is possible to reach the true ground state via QA if we have some preknowledge of the model and are able to prepare the initial ground state in the right topological sector. Depending on the number of sectors $\mathfrak{g}$, at most $\mathfrak{g}$ times can be tried. This is generally hard because even though we know the right sector, it does not mean the corresponding ground state can be efficiently prepared.
On the contrary, QITE algorithm and its NISQ approximate implementation can both successfully address TSO problems demonstrated in our paper. Not only QITE has quadratic advantage on time complexity over QA, but its mechanism which utilizes the property of quantum superposition prevents it from being trapped in some local minima. Topological obstructions have no effects on QITE and thus QITE is a better algorithm for tackling TSO problem than QA. We have to emphasize that the two TSO problems we consider in this work do not even introduce external magnetic fields and it is yet difficult for QA and SQA. If some small random external fields are added, the energy landscape in each sector will be much more glassy and the problems will be more difficult. Unquestionably, QITE can still find the solution as we discussed. The fate of VQITE and VQE in this case is an interesting problem to explore in the future.
@@ Acknowledgements.
We wish to thank Xiaopeng Li and Shuai Yin for helpful discussions. Y.M.D. and Z.Y. thank the start-up fund of Westlake University. The authors acknowledge Beijng PARATERA Tech Co.,Ltd. (https://www.paratera.com/) for providing HPC resources that have contributed to the research results reported within this paper. Y.C.W. acknowledges support from Zhejiang Provincial Natural Science Foundation of China (Grant Nos. LZ23A040003). S.X.Z. is supported by a startup grant in IOP-CAS.
Appendix A Simulation of SQA
@@
For $\hat{H}_{1}$ in the QA Hamiltonian (1), QA always applys uniform magnetic fields on all sites. Different from QA, the SQA method introduces virtual edges to try overcoming the topological obstructions, where a virtual edge along a specified direction will be created imposing strong magnetic fields on some spins. After the spins are polarized, the virtual edge will be glued by decreasing the intensity of the magnetic fields [11].
@@
Here we take the $\hat{H}_{\text{sq}}$ on a $4\times 4$ lattice as an example. Along the vertical direction, there are four virtual edges to be created (See Fig. 11). Here we let $s=t/T$ for convenience. We use $\mathfrak{h}_{j}(s)$ to denote the intensities of magnetic fields acting on the $j$th virtual edge and $\overline{\mathfrak{h}}_{j}(s)$ to denote intensities on those sites that are not included in the $j$th edge at time $s$. Furthermore, we define $s_{\text{max}}<1$ as the time at which we consider all virtual edges have been glued well and $h_{\text{max}}>1$ is the maximum value of $\mathfrak{h}_{j}(s)$.
@@
As shown in Fig. 11, setting $\mathfrak{h}_{1}(0)=h_{\text{max}}$ signifies the creation of the first virtual edge at the beginning of the SQA process. The gluing process of the first edge corresponds to $\mathfrak{h}_{1}(s)=-4h_{\text{max}}s/s_{\text{max}}+h_{\text{max}}$ when $\overline{\mathfrak{h}}_{1}(s)=1-s$, where $s\in[0,(1-h_{\text{max}})/(1-4h_{\text{max}}/s_{\text{max}})]$. Then, a standard QA will be appiled to all sites until $s=s_{\text{max}}/4$. After that, we moves to the second virtual edge by increasing $\mathfrak{h}_{2}(s_{\text{max}}/4)$ from $41-s_{\text{max}}/$ 4 to $h_{\text{max}}$. It is important note that the standard QA is paused until $\mathfrak{h}_{2}(s_{\text{max}})=h_{\text{max}}$. Subsequent to this, we similarly glue the second edge akin to what we do to the first edge. When $s=-4h_{\text{max}}s/s_{\text{max}}+4h_{\text{max}}$, a standard QA will be performed until $s=1$.
@@
Taking into account the time $T_{\text{sweeping}}$ for sequentially creating the virtual edges, the total evolution time should be $T=T_{\text{evo}}+T_{\text{sweeping}}$, where $T_{\text{evo}}$ stands for the time when $s$ is changed. In our simulations for both $\hat{H}_{\text{tri}}$ and $\hat{H}_{\text{sq}}$, we take $h_{\text{max}}=2$ and $s_{\text{max}}=0.8$, which are same as that in Ref. [11]. Besides, the choice of $\Delta t$ and $\Delta s$ follows Appx. II.2.
@@
Appendix B Details of VQITE, Diag-VQITE and VQE
@@
which is exactly the $\Psi^{+}$ in (11), where $H$ denotes Hadamard gate. The variational parameters of the PQC are initially set to follow Gaussian distribution ($\mu=0,\sigma=0.05$). For updating the parameters, we add a small quantity $\epsilon=10^{-4}$ to all the digonal elements $A_{jj}$ in Eq. (15) for the purpose of regularization when solving the linear equations in VQITE and Diag-VQITE. In VQE, we use the stochastic gradient descent to minimize $\langle\phi(\vec{\theta})|\hat{H}|\phi(\vec{\theta})\rangle$.